OpenStudy (anonymous):

find the absolute maximum and minimum values of f(x) = x^3 -4x^2 +7x /x on the interval (0,3] please help on this one?

OpenStudy (anonymous):

Well first, do you see any way to simplify the function?

OpenStudy (anonymous):

find the diferential?

OpenStudy (anonymous):

@jollyjolly0

OpenStudy (anonymous):

is this your function? \[f(x) = \frac{ x^3 -4x^2 + 7x}{ x }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

You should be able simplify that division. Divide everything by x.

OpenStudy (anonymous):

x(x^2 -4x+7)

OpenStudy (anonymous):

But that entire thing is divided by x. So the x you just factored out cancels.

OpenStudy (anonymous):

ok so I am left with x^2-4x+7 ?

OpenStudy (anonymous):

sourwing? please help me?

OpenStudy (anonymous):

then what u wanna do is take the derivative of that and find where the slope would be zero (horizontal) so that u can see where the maxes and mins are from that interval.

OpenStudy (anonymous):

Sorry, I went to eat. Basically, that is a parabola right? Parabolas with a positive second degree term open upwards, so the minimum value is at the vertex. The function increases symmetrically around the vertex, so whichever interval endpoint is furthest from the vertex is the maximum value of the function.

OpenStudy (anonymous):

0 technically isn't included in the interval, so watch out.

OpenStudy (anonymous):

f(x)=x^2-4x+7 f'(x)=2x-4

OpenStudy (anonymous):

2x-4=0

OpenStudy (anonymous):

2x=4 x=2

OpenStudy (anonymous):

so is that my mimimum value?

OpenStudy (loser66):

since you have only 1 critical point, you have to take second derivative to know what it is f"(x) = 2 > 0 always ---> concave up in any interval--> x =2 is abs minimum; no abs max; no relative max/min

OpenStudy (anonymous):

@Loser66 I forgot to put the multiple choices up. they gave me Max: none; min: 4 Max: 7; min: 4 Max: 7; min: 3 Max: none; min: 3 None of these

OpenStudy (anonymous):

so i should choose none of these?

OpenStudy (anonymous):

@phi

OpenStudy (phi):

They want the absolute maximum and minimum values of f(x) = x^3 -4x^2 +7x /x on the interval (0,3] which simplifies to f(x)= x^2 -4x + 7 (x≠0) you found a min at x=2, where f(2)= 3 you should also check the value of f(x) at the boundaries to find the max value *in this interval*

OpenStudy (anonymous):

right so that answwer isn't there

OpenStudy (anonymous):

there is no answer with min: 2

OpenStudy (anonymous):

so i choose "none of these" right?

OpenStudy (phi):

the minimum value of f(x) occurs at x=2. the minimum value itself is 3

OpenStudy (anonymous):

so I need to find the max and its either 7 or 0

OpenStudy (anonymous):

if f(0) is undefined that means it has to be 7 right?

OpenStudy (phi):

You seem to be confusing the x value with the value of f(x) the max or min is f(x) f(0) is undefined, but we can use the limit of f(x) as x->0 In practice, use x=0 in the equation so you want to evaluate f(0) and f(3) one or the other will be the maximum value. A picture is worth a thousand words... but calculus gives you the same answer.

OpenStudy (anonymous):

I have a lot of trouble with the calculus part but the graph was perfect. thank you

OpenStudy (phi):

I went and checked, and technically this function does NOT have a maximum. The idea is that f(x) never reaches the value of 7... there are an infinite number of values that get closer and closer to 7, but never reach it. so the correct answer is no maximum, a minimum of 3

OpenStudy (anonymous):

wow ok. how did you check?

OpenStudy (phi):

I went googling to double check.... I half-remembered that the open interval caused some kind of problem... I found this http://answers.yahoo.com/question/index?qid=20110325091909AA0SPcF

OpenStudy (anonymous):

thank you. thank you sooo much

OpenStudy (anonymous):

@phi could you give me a quick reminder on how to find the concavity of a curve at a point ?

OpenStudy (phi):

I remember the test by remembering that y= x^2 is a "smile" and has a minimum y' = 2x gives the slope at any point x y'' = 2 (2nd derivative) is positive. so y''>0 means MIN y'' <0 means MAX y'' = 0 might mean INFLECTION POINT

OpenStudy (phi):

or see this http://www.math.hmc.edu/calculus/tutorials/secondderiv/

OpenStudy (phi):

Or see Khan's videos (they tend to be short) Here is the first in the section on critical points. http://www.khanacademy.org/math/calculus/derivative_applications/critical_points_graphing/v/minima--maxima-and-critical-points I would watch that one and maybe a few of the others.

OpenStudy (anonymous):

ok thank you very very much